OpenStudy (anonymous):

h4P0 how to solve this?

OpenStudy (anonymous):

i got 2 answers, i can show you how, and you help me with the other 2? :D

OpenStudy (anonymous):

just send him procedure!

OpenStudy (anonymous):

I can't find the resistances

OpenStudy (anonymous):

just apply ohm law and get resistances

OpenStudy (anonymous):

@elenm apply ohm law on it

OpenStudy (anonymous):

hmmm ohms law ? serious? om doing Ri for Vpk of 4v = Ip/Vpk and i got the red cross, thats why i post for help :) do i have something wrong?

OpenStudy (anonymous):

but i get green signal:)by doing this! R=V/I

OpenStudy (anonymous):

@mrvamp CHECK again by applying just ohm law on it

OpenStudy (anonymous):

R=1/(dI/Dv)

OpenStudy (anonymous):

no man sorry i wrote it wron but i did it like that and i got X signal :(

OpenStudy (anonymous):

@mrvamp check that

OpenStudy (anonymous):

d=derivative

OpenStudy (anonymous):

dI=dIp dV=dVpk

OpenStudy (anonymous):

derivative... im stuck i know its an easy one but dont know really, can you give me a "little explain of how to get the apropiate derivative" ??

OpenStudy (anonymous):

apply product rule of derivative by taking p=constant and u know this is exact procedure just try it!

OpenStudy (anonymous):

@elenm u apply that procedure!i think so helpful!

OpenStudy (anonymous):

:( i know it is, issue is im stuck on how to apply

OpenStudy (anonymous):

any book page to refer?

OpenStudy (anonymous):

i think i kind of got it now how but have to read a little more to get it ok, thanks Ali

OpenStudy (anonymous):

@mrvamp its fine !

OpenStudy (anonymous):

I still dont get it, how do I solve for the resistors?

OpenStudy (anonymous):

for the resister you have to solve the derivative.... R=1/(dIp/dVpk) = 1/(d(pv^33/2)/dv = 1/ .003sqrt(v)/1 ....using the V they give you you can solve for R